OpenStudy (anonymous):

How many solutions does the equation have? 4x+2(x-3)=8x+12 A.one solution B.infinite solutions C.no solution

OpenStudy (chosenmatt):

C

OpenStudy (anonymous):

@shinalcantara May you help me with this one ?

OpenStudy (shinalcantara):

4x +2(x-3) = 8x + 12 4x + 2x - 6 = 8x + 12 6x - 6 = 8x + 12 6x - 8x = 12 + 6 -2x = 18 x = -9 then it has one solution